OpenStudy (dragons1415):

Algebra 2 shenanigans >.> Solve the quadratic equation by completing the square. x^2 - 2x = -2

OpenStudy (dragons1415):

Options: A. {1 + i, -1 - i} B. {-1 + i, 1 - i} C. {-1 + i, -1 - i} D. {1 + i, 1 - i}

OpenStudy (dragons1415):

Hate to admit it, but I have absolutely no clue how to do this >.>

OpenStudy (dtan5457):

Are those suppose to be x instead of i's..

OpenStudy (dtan5457):

And I don't see why you would need to complete the square here. x^2-2x=-2 If you add the two x^2-2x+2=0 that's a factorable trinomial

OpenStudy (dragons1415):

The x's are x's... and I just realized what I was doing wrong... I think >.> Might need help later down the line

OpenStudy (dtan5457):

I mean the answers for your option choices..i don't see imaginary solutions

OpenStudy (dragons1415):

Oh no, the i's are i's, which is why I'm getting confused... the teacher might have made a mistake but who knows >.>

OpenStudy (dtan5457):

Oh man..I just said x^2-2x+2=0 is factorable..I must be going crazy.

OpenStudy (dtan5457):

Give me a sec

OpenStudy (dtan5457):

You DO need to complete the square lol

OpenStudy (dragons1415):

lol yeah because the factors of 2 will add up to three :P in that case, I got this: x^2 -2x - 1 = -2 -1

OpenStudy (dragons1415):

which ends up being (x-1)^2 = -3.... and then square root... Dang it, I just needed to think about it more xD Over-complicating things is my specialty smh

OpenStudy (dragons1415):

nope, nevermind... did something wrong again >.> homework + late at night = not a good combo

OpenStudy (dtan5457):

I got it lol x^2-2x=-2 Get a new c value using (b/2)^2 (-2/2)^2=1 add one to both sides x^2-2x+1=-1 factor trinomial (x-1)^2=-1 take square root of both sides x-1=square root of -1 -1=i x-1=i add the one x=1 +/- i

OpenStudy (dtan5457):

i actually forgot how to complete the square for a second

OpenStudy (dtan5457):

happens to everyone (i hope)

OpenStudy (dragons1415):

Thanks sooo much! subtracted a one instead of adding >.> oops

OpenStudy (dragons1415):

Heck, we learned about completing the square days ago and I forgot so I assuming it happens to everyone :P

OpenStudy (dtan5457):

your welcome, to be sure, what is your answer choice?

OpenStudy (dragons1415):

D .-.

OpenStudy (dtan5457):

Yes xD

OpenStudy (dtan5457):

you actually tried and did a lot on this problem, i like that.

OpenStudy (dtan5457):

alot of people just ask for answers

OpenStudy (dtan5457):

thanks for not being one of them

OpenStudy (dragons1415):

You're welcome .-. besides, how will I learn anything if I didn't at least attempt it myself? xD I have to catch mistakes I make so that I (hopefully) don't repeat them .-.

OpenStudy (dtan5457):

i wish everyone had your mentality. i got into plenty of arguments with people too lazy to try anything. anyways, feel free to tag me onto anything, i'll gladly help you if i can.

OpenStudy (dragons1415):

Haha thank you so much! :) I know I said it before, but I want to say it again xD You really helped me out a lot :)